# Limits on planetary-mass primordial black holes from the OGLE high-cadence survey of the Magellanic Clouds

### v1

@@ Abstract
Observations of the Galactic bulge revealed an excess of short-timescale gravitational microlensing events that are generally attributed to a large population of free-floating or wide-orbit exoplanets. However, in recent years, some authors suggested that planetary-mass primordial black holes (PBHs) comprising a substantial fraction (1–10%) of the dark matter in the Milky Way may be responsible for these events. If that was the case, a large number of short-timescale microlensing events should also be seen toward the Magellanic Clouds. Here we report the results of a high-cadence survey of the Magellanic Clouds carried out from October 2022 through May 2024 as part of the Optical Gravitational Lensing Experiment (OGLE). We observed almost 35 million source stars located in the central regions of the Large and Small Magellanic Clouds and found only one long-timescale microlensing event candidate. No short-timescale events were detected despite high sensitivity to such events. That allows us to infer the strongest available limits on the frequency of planetary-mass PBHs in dark matter. We find that PBHs and other compact objects with masses from $1.4\times 10^{-8}\,M_{\odot}$ (half of the Moon mass) to $0.013\,M_{\odot}$ (planet/brown dwarf boundary) may comprise at most 1% of dark matter. That rules out the PBH origin hypothesis for the short-timescale events detected toward the Galactic bulge and indicates they are caused by the population of free-floating or wide-orbit planets.
@@ 1 Introduction
Figure 1: 95% upper limits on the fraction of dark matter in the form of PBHs and other compact objects. The hatched area marks the region of the PBH parameter space obtained by Niikura et al. (2019a), under assumption that short-timescale microlensing events detected by Mróz et al. (2017) in the OGLE-IV Galactic bulge high-cadence survey are due to planetary-mass PBHs. The shaded purple region corresponds to the 95% confidence limit on $f$ found in this work. Other limits are adopted from Tisserand et al. (2007) (EROS), Wyrzykowski et al. (2011) (OGLE-III), Griest et al. (2013) (Kepler), Niikura et al. (2019b) (HSC), Blaineau et al. (2022) (MACHO+EROS), and Mróz et al. (2024b) (OGLE-III+OGLE-IV).
Soon after the discovery of six short-timescale microlensing events by Mróz et al. (2017) in the data from the high-cadence Galactic bulge survey by the fourth phase of the Optical Gravitational Lensing Experiment (OGLE), the paper by Niikura et al. (2019a) raised speculations about their origin. These authors put forward a claim that the short-timescale events discovered by OGLE were caused by planetary-mass primordial black holes (PBHs), not free-floating or wide-orbit planets, as argued in the original discovery paper. Niikura et al. (2019a) found that if the lenses were indeed PBHs and if their distribution in the Milky Way followed the Navarro-Frenk-White model (Navarro et al., 1997; Klypin et al., 2002), then PBHs of a few $M_{\oplus}$ could comprise $1\!-\!10\%$ of the dark matter in the Milky Way (see the hatched area in Figure 1).
These claims inspired many theoretical follow-up studies, which attempted to explain the origin of planetary-mass PBHs allegedly detected in the OGLE data. For example, a purported population of planetary-mass PBHs was claimed to be consistent with different models of inflation (e.g., Tada & Yokoyama, 2019; Fu et al., 2019; Motohashi et al., 2020; Teimoori et al., 2021a, b; Solbi & Karami, 2021; Heydari & Karami, 2022; Zhang, 2022; Yi, 2023; Fu & Chen, 2023; Fu & Wang, 2023; Flores & Kusenko, 2023; Ashrafzadeh & Karami, 2024; Heydari & Karami, 2024a, b; Yi et al., 2024; Cai et al., 2024; Yang et al., 2024). Several authors (e.g., Domènech & Pi, 2022; Zhang, 2022; Yi, 2023; Inomata et al., 2024; Domènech et al., 2024; Yang et al., 2024; Yi et al., 2024) proposed that planetary-mass PBHs may be linked to the detection of the stochastic gravitational wave background (Arzoumanian et al., 2020; Agazie et al., 2023a, b; EPTA Collaboration et al., 2023a, b; Reardon et al., 2023; Zic et al., 2023; Xu et al., 2023), as they both can have a common origin in the primordial curvature perturbations.
@@
We expect that planets (either free-floating or bound to their host stars) should follow the distribution of stars in the Milky Way, whereas PBHs would track the distribution of dark matter. Distinguishing between the free-floating/wide-orbit planet and the PBH hypotheses for Galactic bulge microlensing events would require precise measurements of the distances to lensing objects. [^1] The populations of FFPs and PBHs can, in theory, also be separated in a statistical sense based on $t_{\rm E}$ or $\theta_{\rm E}$ measurements for a large sample of events (e.g., DeRocco et al., 2024). That would, in turn, require measurements of microlensing parallaxes of short-timescale events, which are, in practice, nearly impossible with the current technology. (Such measurements may be feasible in the future with simultaneous observations from two separated observatories, e.g., Zhu & Gould 2016; Bachelet & Penny 2019; Ban 2020; Gould et al. 2021; Yan & Zhu 2022.)
@@ 2 Data
The data analyzed in this study were collected as part of the OGLE-IV survey (Udalski et al., 2015). The survey uses the dedicated 1.3-m Warsaw Telescope located at Las Campanas Observatory, Chile. (The observatory is operated by the Carnegie Institution for Science.) The telescope is equipped with a mosaic CCD camera comprising thirty-two $2048\times 4102$ detectors. The pixel scale is 0.26 arcsec pixel-1, providing a total field of view of 1.4 deg2. Udalski et al. (2015) presented a detailed description of this instrument.
The Magellanic Clouds have been a target of the OGLE survey since its second phase (Udalski et al., 1997; Udalski, 2003; Udalski et al., 2015). However, they were observed at a moderate cadence, typically 1–5 days. When OGLE-IV operations were resumed after the COVID-19 pandemic in August 2022, we decided to observe some Magellanic Clouds fields at high cadence. We selected four fields in the Small Magellanic Cloud (SMC) and five in the LMC, which contained the largest number of stars. They cover an area of approximately 12.4 deg2 and contain almost 35 million source star brighter than $I=22$. Basic information about these fields (equatorial coordinates, total number of epochs in 2022/2023 and 2023/2024 observing seasons, and the number of microlensing source stars) is presented in Table 1. The location of the fields in the sky is shown in Figure 2.
The four SMC fields were observed up to 21 times per night (with a median of 9 times per night) with a median cadence of 16 minutes. The median cadence in the LMC was slightly longer (20 minutes). The LMC fields were observed up to 23 times per night (with a median of 12 times per night). High-cadence survey observations of the SMC were conducted from 2022 Oct 9 to 2022 Dec 11 and from 2023 Aug 11 to 2024 Jan 29 (in total, 236 days). The LMC was observed at high cadence from 2022 Oct 4 to 2023 Apr 30 and from 2023 Sep 15 to 2024 Apr 27 (in total, 435 days).
@@
Figure 2: Fields of the OGLE High-Cadence Magellanic Clouds Survey; left panel, LMC; right panel, SMC. The background images of the LMC and SMC were generated with bsrender written by Kevin Loch, using the ESA/Gaia database.
@@ 3 Search for Events
The methods that we employed for searching for microlensing events in the time-series data are necessarily similar to those used in our previous works (Mróz et al., 2017, 2024a). One important difference arises from the fact that we expect that some light curves of microlensing events due to planetary-mass objects may be affected by finite-source effects. The selection of events is carried out in three steps. First, we search for objects exhibiting any brightening with respect to the flat light curve. Then, we remove obvious non-microlensing light curves, and finally, check whether a microlensing model can fit the light curve well. The selection criteria are devised to maximize the detection efficiency (see Section 5) while keeping the contamination from non-microlensing light curves as small as possible. All selection cuts are summarized in Table 2.
@@ 3.1 Cut 1
We searched for microlensing events using only the data collected between 2022 Aug 9 and 2024 May 20. The error bars reported by the photometric pipeline were rescaled following the procedure described by Skowron et al. (2016) and Mróz et al. (2024a), and we converted magnitudes into flux units. Then, for each light curve containing at least 30 data points, we placed a window lasting 120 days and calculated the mean flux $F_{\rm base}$ and its standard deviation $\sigma_{\rm base}$ using data points outside the window (after removing 4-sigma outliers). We searched for at least $n_{\rm bump}=5$ consecutive data points within the window, whose flux was greater than $F_{\rm base}+3\sigma_{\rm base}$. We called such data points a “bump”. In case such a bump was identified in the light curve, we calculated the quantity $\chi_{3+}=\sum_{i}(F_{i}-F_{\rm base})/\sigma_{i}$, where $(F_{i},\sigma_{i})$ is the flux and its uncertainty of the $i$th data point within he bump. Subsequently, the window was shifted by 40 days, and the whole procedure was repeated until the end of the light curve was reached. From all iterations of the algorithm, we selected a bump with the largest value of $\chi_{3+}$ and required $\chi_{3+}\geq 40$.
In addition, we required that during at least $n_{\rm DIA}=3$ epochs the magnified source was detected by the DIA pipeline on subtracted images (c.f., Mróz et al. 2024a), and the amplitude of the bump should be $\Delta m\geq 0.05$ mag. The latter cut was devised to remove contamination from OGLE small amplitude red giants (Wray et al., 2004), whose low-amplitude pulsations were picked up by our algorithm as candidate microlensing events. We also required that no significant variability was detected outside the magnified part of the light curve. To quantify that, we calculated $\chi^{2}_{\rm out}=\sum_{i}(F_{i}-F_{\rm base})^{2}/\sigma_{i}^{2}$, where the summation was performed over all data points outside the window, and required $\chi^{2}_{\rm out}/\mathrm{d.o.f.}\leq 2$. In this step, we reduced the number of analyzed light curves from over 17.6 million to 2538.
@@ 3.2 Cut 2
In the second step, we identified false positives and removed them from the sample. They could be classified into three main (non-exclusive) classes: (1) stars with multiple bumps (brightenings) in their light curves, (2) stars located in the “blue-bumper” region in the color–magnitude diagram (CMD), and (3) photometry artifacts.
First, we required that the full OGLE-III (if available) and OGLE-IV light curve contains only one bump. To verify that, we algorithmically removed magnified data points from the main brightening and ran the event-finding algorithm described above using the remaining data. If the algorithm detected another bump with $\chi_{3+}\geq 30$ and $n_{\rm bump}\geq 5$ (note these conditions are more lax than those imposed on the primary bump), we discarded such the light curve.
We then obtained de-reddened colors $(V-I)_{0}$ and extinction-corrected mean magnitudes $I_{0}$ of the selected objects using reddening maps by Skowron et al. (2021). For the LMC fields, we removed stars located in the region of the CMD delimited by $(V-I)_{0}\leq 0.5$ and $I_{0}\leq 19.5$. This CMD region contains stars called “blue-bumpers”, exhibiting repeating outbursts that sometimes can be mistaken with microlensing events (c.f., Alcock et al., 1997; Mróz et al., 2024a). For the SMC fields, the brightness limit was slightly fainter, $I_{0}\leq 20.0$, as the SMC distance modulus is 0.5 mag larger than that of the LMC.
When analyzing selected light curves, we noticed that some apparent bumps were present during the same night in the light curves of many stars located in the same field or on the same CCD detector. These were spurious non-astrophysical signals produced by artifacts that were present in the reduced CCD images. They may have been caused by reflections of the light within the telescope optics or flawed calibration images (flats). We thus discarded stars that showed bumps during a few “bad” nights. After these steps, we were left with 308 light curves.
@@ 3.3 Cut 3
In the last step, we fitted microlensing models to the light curves of the remaining objects and selected brightenings that were consistent with microlensing. We considered three types of models. In the standard, point-source point-lens (PSPL) model, the magnification is given by the formula:
@@
where $F_{0}$ is the baseline flux and $f_{\rm s}$ is the dimensionless blending parameter. We considered two cases. If we fix $f_{\rm s}=1$ (four-parameter model), we assume that the whole observed flux comes from the source star. If we allow $f_{\rm s}\neq 1$ (five-parameter model), then we take into account the possibility of blended light which is not magnified during the event and may come from the lens itself or unrelated stars within the seeing disk of the event.
@@
using the GNU Scientific Library [^2] https://www.gnu.org/software/gsl/ implementation of the Levenberg–Marquardt algorithm (Levenberg, 1944; Marquardt, 1963). The algorithm returned not only the best-fit values, but also their covariance matrix. We fitted three types of models: four- and five-parameter PSPL and five-parameter ESPL. However, we adopted only one of them as the “fiducial” model, using the following algorithm.
If the finite-source effects are present in the light curve and the ESPL fit results in a significantly smaller $\chi^{2}$ than the PSPL fits, we adopted the ESPL fit as fiducial. Specifically, we required that $\Delta\chi^{2}\equiv\chi^{2}_{\rm PSPL}-\chi^{2}_{\rm ESPL}\geq 16$ (we checked that changing this parameter within a reasonable range did not affect our results). If $\Delta\chi^{2}<16$ and the five-parameter PSPL fit did not converge or it converged to an unphysical value of the blending parameter ($f_{\rm s}-\sigma(f_{\rm s})\geq 1$), then we used the four-parameter PSPL fit instead. Otherwise, we adopted the five-parameter PSPL fit.
@@
All criteria were met by only two objects, which we describe in more detail below. We treat one of them as a microlensing event candidate and call it OGLE-LMC-20. Furthermore, we argue that the second object is most likely a stellar flare. In addition to that, we visually inspected the light curves of all 2538 objects that met the “cut 1” criteria. However, we did not find any additional candidates for microlensing events.

### v2

@@ Abstract
Observations of the Galactic bulge revealed an excess of short-timescale gravitational microlensing events that are generally attributed to a large population of free-floating or wide-orbit exoplanets. However, in recent years, some authors suggested that planetary-mass primordial black holes (PBHs) comprising a substantial fraction (1%–10%) of the dark matter in the Milky Way may be responsible for these events. If that was the case, a large number of short-timescale microlensing events should also be seen toward the Magellanic Clouds. Here, we report the results of a high-cadence survey of the Magellanic Clouds carried out from 2022 October through 2024 May as part of the Optical Gravitational Lensing Experiment. We observed almost 35 million source stars located in the central regions of the Large and Small Magellanic Clouds and found only one long-timescale microlensing event candidate. No short-timescale events were detected despite high sensitivity to such events. That allows us to infer the strongest available limits on the frequency of planetary-mass PBHs in dark matter. We find that PBHs and other compact objects with masses from $1.4\times 10^{-8}\,M_{\odot}$ (half of the Moon mass) to $0.013\,M_{\odot}$ (planet/brown dwarf boundary) may comprise at most 1% of dark matter. That rules out the PBH origin hypothesis for the short-timescale events detected toward the Galactic bulge and indicates they are caused by the population of free-floating or wide-orbit planets.
@@ 1 Introduction
Figure 1: The 95% upper limits on the fraction of dark matter in the form of PBHs and other compact objects. The hatched area marks the region of the PBH parameter space obtained by Niikura et al. (2019a), under assumption that short-timescale microlensing events detected by Mróz et al. (2017) in the OGLE-IV Galactic bulge high-cadence survey are due to planetary-mass PBHs. The shaded purple region corresponds to the 95% confidence limit on $f$ found in this work. Other limits are adopted from Tisserand et al. (2007) (EROS), Wyrzykowski et al. (2011) (OGLE-III), Griest et al. (2013) (Kepler), Niikura et al. (2019b) (HSC), Blaineau et al. (2022) (MACHO+EROS), and Mróz et al. (2024b) (OGLE-III+OGLE-IV).
Soon after the discovery of six short-timescale microlensing events by Mróz et al. (2017) in the data from the high-cadence Galactic bulge survey by the fourth phase of the Optical Gravitational Lensing Experiment (OGLE), the paper by Niikura et al. (2019a) raised speculations about their origin. These authors put forward a claim that the short-timescale events discovered by OGLE were caused by planetary-mass primordial black holes (PBHs), not free-floating or wide-orbit planets, as argued in the original discovery paper. Niikura et al. (2019a) found that if the lenses were indeed PBHs and if their distribution in the Milky Way followed the Navarro–Frenk–White model (Navarro et al., 1997; Klypin et al., 2002), then PBHs of a few $M_{\oplus}$ could comprise $1\%\!-\!10\%$ of the dark matter in the Milky Way (see the hatched area in Figure 1).
These claims inspired many theoretical follow-up studies, which attempted to explain the origin of planetary-mass PBHs allegedly detected in the OGLE data. For example, a purported population of planetary-mass PBHs was claimed to be consistent with different models of inflation (e.g., Tada & Yokoyama, 2019; Fu et al., 2019; Motohashi et al., 2020; Teimoori et al., 2021a, b; Solbi & Karami, 2021; Heydari & Karami, 2022; Zhang, 2022; Yi, 2023; Fu & Chen, 2023; Fu & Wang, 2023; Flores & Kusenko, 2023; Ashrafzadeh & Karami, 2024; Heydari & Karami, 2024a, b; Yi et al., 2024; Cai et al., 2024; Yang et al., 2024). Several authors (e.g., Domènech & Pi, 2022; Zhang, 2022; Yi, 2023; Franciolini et al., 2023; Inomata et al., 2024; Domènech et al., 2024; Yang et al., 2024; Yi et al., 2024) proposed that planetary-mass PBHs may be linked to the detection of the stochastic gravitational wave background (Arzoumanian et al., 2020; Agazie et al., 2023a, b; EPTA Collaboration et al., 2023a, b; Reardon et al., 2023; Zic et al., 2023; Xu et al., 2023), as they both can have a common origin in the primordial curvature perturbations.
@@
We expect that planets (either free-floating or bound to their host stars) should follow the distribution of stars in the Milky Way, whereas PBHs would track the distribution of dark matter. Distinguishing between the free-floating/wide-orbit planet and the PBH hypotheses for Galactic bulge microlensing events would require precise measurements of the distances to lensing objects. [^1] The populations of FFPs and PBHs can, in theory, also be separated in a statistical sense based on $t_{\rm E}$ or $\theta_{\rm E}$ measurements for a large sample of events (e.g., DeRocco et al., 2024). That would, in turn, require measurements of microlensing parallaxes of short-timescale events, which are, in practice, nearly impossible with the current technology. (Such measurements may be feasible in the future with simultaneous observations from two separated observatories; e.g., Zhu & Gould 2016; Bachelet & Penny 2019; Ban 2020; Gould et al. 2021; Yan & Zhu 2022.)
@@ 2 Data
The data analyzed in this study were collected as part of the OGLE-IV survey (Udalski et al., 2015). The survey uses the dedicated 1.3 m Warsaw Telescope located at Las Campanas Observatory, Chile. (The observatory is operated by the Carnegie Institution for Science.) The telescope is equipped with a mosaic CCD camera comprising thirty-two $2048\times 4102$ detectors. The pixel scale is 0.26 arcsec pixel-1, providing a total field of view of 1.4 deg2. Udalski et al. (2015) presented a detailed description of this instrument.
The Magellanic Clouds have been a target of the OGLE survey since its second phase (Udalski et al., 1997; Udalski, 2003; Udalski et al., 2015). However, they were observed at a moderate cadence, typically 1–5 days. When OGLE-IV operations were resumed after the COVID-19 pandemic in 2022 August, we decided to observe some Magellanic Clouds fields at high cadence. We selected four fields in the Small Magellanic Cloud (SMC) and five in the LMC, which contained the largest number of stars. They cover an area of approximately 12.4 deg2 and contain almost 35 million source stars brighter than $I=22$. Basic information about these fields (equatorial coordinates, total number of epochs in 2022/2023 and 2023/2024 observing seasons, and the number of microlensing source stars) is presented in Table 1. The location of the fields in the sky is shown in Figure 2.
The four SMC fields were observed up to 21 times per night (with a median of 9 times per night) with a median cadence of 16 minutes. The median cadence in the LMC was slightly longer (20 minutes). The LMC fields were observed up to 23 times per night (with a median of 12 times per night). High-cadence survey observations of the SMC were conducted from 2022 October 9 to 2022 December 11 and from 2023 August 11 to 2024 January 29 (in total, 236 days). The LMC was observed at high cadence from 2022 October 4 to 2023 April 30 and from 2023 September 15 to 2024 April 27 (in total, 435 days).
@@
Figure 2: Fields of the OGLE High-cadence Magellanic Clouds Survey; left panel—LMC; right panel—SMC. The background images of the LMC and SMC were generated with bsrender written by Kevin Loch, using the ESA/Gaia database.
@@ 3 Search for Events
The methods that we employed for searching for microlensing events in the time-series data are necessarily similar to those used in our previous works (Mróz et al., 2017, 2024a). One important difference arises from the fact that we expect that some light curves of microlensing events due to planetary-mass objects may be affected by finite-source effects. The selection of events is carried out in three steps. First, we search for objects exhibiting any brightening with respect to the flat light curve. Then, we remove obvious nonmicrolensing light curves and, finally, check whether a microlensing model can fit the light curve well. The selection criteria are devised to maximize the detection efficiency (see Section 5) while keeping the contamination from nonmicrolensing light curves as small as possible. All selection cuts are summarized in Table 2.
@@ 3.1 Cut 1
We searched for microlensing events using only the data collected between 2022 August 9 and 2024 May 20. The error bars reported by the photometric pipeline were rescaled following the procedure described by Skowron et al. (2016) and Mróz et al. (2024a), and we converted magnitudes into flux units. Then, for each light curve containing at least 30 data points, we placed a window lasting 120 days and calculated the mean flux $F_{\rm base}$ and its standard deviation $\sigma_{\rm base}$ using data points outside the window (after removing $4\sigma$ outliers). We searched for at least $n_{\rm bump}=5$ consecutive data points within the window, whose flux was greater than $F_{\rm base}+3\sigma_{\rm base}$. We called such data points a “bump.” In case such a bump was identified in the light curve, we calculated the quantity $\chi_{3+}=\sum_{i}(F_{i}-F_{\rm base})/\sigma_{i}$, where $(F_{i},\sigma_{i})$ is the flux and its uncertainty of the $i$th data point within the bump. Subsequently, the window was shifted by 40 days, and the whole procedure was repeated until the end of the light curve was reached. From all iterations of the algorithm, we selected a bump with the largest value of $\chi_{3+}$ and required $\chi_{3+}\geq 40$.
In addition, we required that during at least $n_{\rm DIA}=3$ epochs the magnified source was detected by the DIA pipeline on subtracted images (see, Mróz et al. 2024a), and the amplitude of the bump should be $\Delta m\geq 0.05$ mag. The latter cut was devised to remove contamination from OGLE small-amplitude red giants (Wray et al., 2004), whose low-amplitude pulsations were picked up by our algorithm as candidate microlensing events. We also required that no significant variability was detected outside the magnified part of the light curve. To quantify that, we calculated $\chi^{2}_{\rm out}=\sum_{i}(F_{i}-F_{\rm base})^{2}/\sigma_{i}^{2}$, where the summation was performed over all data points outside the window and required $\chi^{2}_{\rm out}/\mathrm{d.o.f.}\leq 2$. In this step, we reduced the number of analyzed light curves from over 17.6 million to 2538.
@@ 3.2 Cut 2
In the second step, we identified false positives and removed them from the sample. They could be classified into three main (nonexclusive) classes: (1) stars with multiple bumps (brightenings) in their light curves, (2) stars located in the “blue-bumper” region in the color–magnitude diagram (CMD), and (3) photometry artifacts.
First, we required that the full OGLE-III (if available) and OGLE-IV light curve contains only one bump. To verify that, we algorithmically removed magnified data points from the main brightening and ran the event-finding algorithm described above using the remaining data. If the algorithm detected another bump with $\chi_{3+}\geq 30$ and $n_{\rm bump}\geq 5$ (note these conditions are more lax than those imposed on the primary bump), we discarded such a light curve.
We then obtained dereddened colors $(V-I)_{0}$ and extinction-corrected mean magnitudes $I_{0}$ of the selected objects using reddening maps by Skowron et al. (2021). For the LMC fields, we removed stars located in the region of the CMD delimited by $(V-I)_{0}\leq 0.5$ and $I_{0}\leq 19.5$. This CMD region contains stars called “blue-bumpers,” exhibiting repeating outbursts that sometimes can be mistaken with microlensing events (see, Alcock et al., 1997; Mróz et al., 2024a). For the SMC fields, the brightness limit was slightly fainter, $I_{0}\leq 20.0$, as the SMC distance modulus is 0.5 mag larger than that of the LMC.
When analyzing selected light curves, we noticed that some apparent bumps were present during the same night in the light curves of many stars located in the same field or on the same CCD detector. These were spurious nonastrophysical signals produced by artifacts that were present in the reduced CCD images. They may have been caused by reflections of the light within the telescope optics or flawed calibration images (flats). We thus discarded stars that showed bumps during a few “bad” nights. After these steps, we were left with 308 light curves.
@@ 3.3 Cut 3
In the last step, we fitted microlensing models to the light curves of the remaining objects and selected brightenings that were consistent with microlensing. We considered three types of models. In the standard, point-source point-lens (PSPL) model, the magnification is given by the following formula:
@@
where $F_{0}$ is the baseline flux and $f_{\rm s}$ is the dimensionless blending parameter. We considered two cases. If we fix $f_{\rm s}=1$ (four-parameter model), we assume that the whole observed flux comes from the source star. If we allow $f_{\rm s}\neq 1$ (five-parameter model), then we take into account the possibility of blended light that is not magnified during the event and may come from the lens itself or unrelated stars within the seeing disk of the event.
@@
using the GNU Scientific Library [^2] https://www.gnu.org/software/gsl/ implementation of the Levenberg–Marquardt algorithm (Levenberg, 1944; Marquardt, 1963). The algorithm returned not only the best-fit values but also their covariance matrix. We fitted three types of models: four- and five-parameter PSPL and five-parameter ESPL. However, we adopted only one of them as the “fiducial” model, using the following algorithm.
If the finite-source effects were present in the light curve and the ESPL fit resulted in a significantly smaller $\chi^{2}$ than the PSPL fits, we adopted the ESPL fit as fiducial. Specifically, we required that $\Delta\chi^{2}\equiv\chi^{2}_{\rm PSPL}-\chi^{2}_{\rm ESPL}\geq 16$ (we checked that changing this parameter within a reasonable range did not affect our results). If $\Delta\chi^{2}<16$ and the five-parameter PSPL fit did not converge or it converged to an unphysical value of the blending parameter ($f_{\rm s}-\sigma(f_{\rm s})\geq 1$), then we used the four-parameter PSPL fit instead. Otherwise, we adopted the five-parameter PSPL fit.
@@
All criteria were met by only two objects, which we describe in more detail below. Their properties are summarized in Table 3. We treat one of them as a microlensing event candidate and call it OGLE-LMC-20. Furthermore, we argue that the second object is most likely a stellar flare. In addition to that, we visually inspected the light curves of all 2538 objects that met the “cut 1” criteria. However, we did not find any additional candidates for microlensing events.